OpenStudy (anonymous):

what is the diffrence between three times a number and 10 is 17

OpenStudy (anonymous):

? = 3x - 3, when 10 = 17. The logic makes no sense, restate your question please?!?!

OpenStudy (alexwee123):

3x-10=17

OpenStudy (anonymous):

ty alexwee 123

OpenStudy (alexwee123):

:)

OpenStudy (anonymous):

lol good thing someone speaks mangled algebra, I was really lost there.. xD

OpenStudy (anonymous):

i got the question out of the actual math book word for word

OpenStudy (anonymous):

(diffrence between three times a number and 10) is (17) would help lol

OpenStudy (anonymous):

ok i will be more specific next time

OpenStudy (anonymous):

3x-13=-21 solove in steps please

OpenStudy (alexwee123):

add 13 on both sides then divide by 3 on both sides

OpenStudy (alexwee123):